Derek has the opportunity to buy a money machine today. The money machine will pay Derek $14,605.00 exactly 4.00 years from today. Assuming that Derek believes the appropriate discount rate is 10.00%, how much is he willing to pay for this money machine?

Suppose you deposit $2,368.00 into an account today that earns 10.00%. In 22.00 years the account will be worth $________.

Suppose you deposit $2,573.00 into an account today. In 8.00 years the account is worth $3,815.00. The account earned ____% per year.
Answer #1

1). PV = FV / (1 + r)n

= $14,605 / 1.104 = $9,975.41

2). FV = PV x (1 + r)n

= $2,368 x 1.1022 = $19,276.17

3). r = [FV / PV]1/n - 1

= [$3,815 / $2,573]1/8 - 1 = 1.0505 - 1 = 0.0505, or 5.05%